OpenStudy (tttruong):

what is the equation of the plane that passes through the point (1, -2, 4) and is perpendicular to the vector -i-3k

OpenStudy (518nad):

you have vector <-1,0,3> as the normal line and a point (1,-2,4) on the plane

OpenStudy (518nad):

a place equation is given by ax+by+cz=d where a,b,c is from the normal line to this plane <a,b,c>

OpenStudy (518nad):

thus we have -1x+0y+3z=d now to find d, plug in your point and see the d that satisfied (1,-2,4) -1*1+0*-2+3*4=d d=10 -1x+0y+3z=10 x+3z=10 is our plane

OpenStudy (518nad):

in the 2nd post that should say, a plane* equation

OpenStudy (518nad):

do you have any questions

OpenStudy (tttruong):

how can you find the vector (-1, 0, 3)

OpenStudy (518nad):

they gave u that

OpenStudy (518nad):

perpendicular to the vector -i-3k

OpenStudy (tttruong):

i dont understand how i can know from the vector they give -i-3k will be (-1, 0,3)

OpenStudy (518nad):



OpenStudy (518nad):

i j k are like x,y,z

OpenStudy (518nad):

i j k are unit vectors along each axis here

OpenStudy (518nad):

so they say u are going -1 on the i direction, and -3 on the k direction, by not specifying what j is in this case, its already taken to be 0

OpenStudy (tttruong):

oh i see it know

OpenStudy (tttruong):

thx so much

OpenStudy (518nad):

i can give u a more dot product way of solving for this, if you a bit unsettled by why the normal line is the coeffecients a,b,c

OpenStudy (518nad):

it is good to understand both

OpenStudy (danjs):

equation for a plane, normal n, and position vector r0 \[\large n*(r-r _{o)}=0\]

OpenStudy (518nad):

yes i was gonna go into that definition, if he wanted

OpenStudy (danjs):

just remember that, need a point on the plane and a normal direction \[<a,b,c>*(<x,y,z>-<x _{o},y _{0},z _{o}>)=0\]

OpenStudy (danjs):

it will expand out to the scalar equation for the plane

OpenStudy (tttruong):

are they the same with parallel line?

OpenStudy (518nad):

if this plane passes through the origin then the vectors on this plane are of the form <x,y,z> and we know dot product of 2 perpendicular vectors give u 0 so <-1,0,-3> dot <x,y,z> = 0 -1x -3z=0 but remember this is going through the origin, so to simply displace this to the point we want (1, -2, 4) -1(x-1) -3(z-4) = 0

OpenStudy (518nad):

if you remember or understand the line equation form y-b= m(x-a) is a line of the slope m at point (a,b)

OpenStudy (518nad):

this is exactly what we are doing to displace this origin passing plane to make the new 'origin' at the point we want

OpenStudy (518nad):

u can think about it as a new set of variables u,v,t, for the x,y,z... i dunno i wont go into all this lol, its something for you to make sense of, not my own way of making it comfortable for myself

OpenStudy (tttruong):

thank you i get it now